OpenStudy (anonymous):

Imaginary Numbers and Conjugates: If z=2-3i and w=1+2i, evaluate z+w (conjugates of z, and w, and of z+w)

OpenStudy (anonymous):

If \[z=2-3i\] and \[w=1+2i\] evaluate 

OpenStudy (anonymous):

The conjugate of the sums of the conjugates?

OpenStudy (anonymous):

I believe, but I do not know for sure?

OpenStudy (anonymous):

\[\begin{split} \overline{\overline{a+bi}+\overline{c+di}} &= \overline{a-bi+c-di} \\ &=\overline{a+c-(b+d)i}\\ &=(a+c)+(b+d)i \end{split}\]

OpenStudy (anonymous):

The conjugate of the sum of the conjugates seems to just be the plain old sum.

OpenStudy (anonymous):

I am going to try it right now, has to be better than what I had come up with.

OpenStudy (anonymous):

\[ z=2−3i \implies \overline z = 2+3i \]

OpenStudy (anonymous):

\[ w=1+2i\implies \overline w=1-2i \]

OpenStudy (anonymous):

\[ \overline z+\overline w=(2+3i)+(1−2i)=3+i \]

OpenStudy (anonymous):

\[ \overline{\overline z+\overline w}=\overline{3+i}=3-i \]

OpenStudy (anonymous):

\[ z+w=(2−3i)+(1+2i)=3-i \]

OpenStudy (anonymous):

Seems consistent.

OpenStudy (anonymous):

Awesome, so seems as though I got half of the answers right, the way I worked it out, I ended up with 4 answers. Mistook what I was looking at. Thank you so much!